You notice two balls rolling down a hill. Both balls were initially at rest, both have the same mass and same radius, but one of them is clearly rolling down faster than the other. What could explain this?
Answer #1

If both balls have same mass and same radius and one of them is rolling faster then this ball must be a solid ball and slower ball is Hollow spherical ball.

acceleration of a rolling ball is a = g sin\Theta /^{\sqrt{1+(K^2/R^2)}}

where k is radius of gyration

for solid sphere K = sqrt(2/5)R

for hollow sphere K = sqrt(2/3)R

so by putting these values we get

for solid ball a = 5/7(g sin theta)

for hollow ball a = 3/5(g sin theta)

so acceleration of solid ball is more compare to hollow ball.